the formula A= 1/2bh gives the area A for a triangle
Solve for H
Solve for B
Find H if B= 10in A= 80in2

Answer by wuwei96815(245)

A = 1/2bh
Multiply both sides by 2
2A = bh
Divide both sides by b
2A/b = h

A = 1/2bh
Multiply both sides by 2
2A = bh
Divide both sides by h
2A/h = b

If b = 10" and A = 80 inches2, what is h?
Use substitution
2A/b = h
2(80)/10 = h
160/10 = h
16" = h